OpenStudy (anonymous):

How much time will a $500 investment need to earn $100 in interest at a rate of 10% per year?

OpenStudy (amistre64):

simple or compounded interest?

OpenStudy (amistre64):

simple is just: A = Prt 100 = 500(.10) t solve for t

OpenStudy (amistre64):

if its compounded once a year ... its a little more complex but: B = B(1+r)^t 600 = 500(1.10)^t 6/5 = 1.10 ^t ln(6/5) = t ln(1.10) ln(6/5)/ln(1.10) = t

OpenStudy (anonymous):

ok thanx so much

OpenStudy (amistre64):

simple interest is 2 years compounded is 1.9 years so its prolly a simple interest calculation

OpenStudy (anonymous):

i think it is

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

thanx